Solve the differential equation


(x² + y²)dx-xydy = 0


1
Expert's answer
2022-01-14T06:09:11-0500

Solution

Given equation may be rewritten in the form

"\frac{dy}{dx}=\frac{x^2+y^2}{xy}=\frac{x}{y}+\frac{y}{x}"

Let z=y/x => "\frac{dy}{dx}=\frac{dz}{dx}x+z"  => "\frac{dz}{dx}x+z\ =\frac{1}{z}+z"  => zdz = dx/x  =>   "\int{zdz=\int\frac{dx}{x}}"    => z2/2=ln|x|+C   =>  C = z2/2 - ln|x| = y2/(2x2)-ln|x|   =>  "y(x)=\sqrt{2x^2\left(C+ln|x|\right)}"

Answer

"y(x)=\sqrt{2x^2\left(C+ln|x|\right)}"